OpenStudy (anonymous):

Can anyone SHOW me how to get the answer to the problem: |9+6x|=4 ?

OpenStudy (anonymous):

So |9+6x| just means absolute value, which just means 9+6x. 9+6x = 4, solve algebraically. Subtract 9 both sides to get 6x alone. 4-9 = -5. 6x = -5 Divide by 6 to get X. x = -5/6

OpenStudy (anonymous):

Absolute value of 9+6x has two solutions. Nielglen covered the obvious solution, where it equals itself. However, it can also be equal to -1 times itself. To find your second solution, solve the equation\[(-1)(9+6x)=4\]

OpenStudy (anonymous):

when x <-9/6 we will open absolute value with a negetive sign that is -9-6x=4 6x=-13 x=-13/6 when x>-9/6 we will open absolute value with a positive sign that is 9+6x=4 5=6x x=5/6

OpenStudy (anonymous):

Thanks Xeph, there are holes in my Algebra. That's why we're all here! Rishabh, you made a mistake in your second part. You subtracted 4 from 9 and got 5 and then simply moved 6x to the right. You need to subtract 6x. Answer is -5/6 = x.